OpenStudy (marziajackson$):

Last Question. What is the value of the expression 1/5^-5 ? A. 1/3125 B. 1/25 C. 25 D. 3125

OpenStudy (marziajackson$):

@Michele_Laino

Nnesha (nnesha):

flip the fraction to change negative to positive exponent \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

Nnesha (nnesha):

\[\huge\rm \frac{ x^{-m} }{ 1 }=\frac{ 1 }{ x^m }\] what is the reciprocal of \[\frac{ 1 }{ 5^{-5} }\] ? do you know ?

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{{{5^{ - 5}}}} = {5^5} = 5 \times 5 \times 5 \times 5 \times 5 = ...?\]

Nnesha (nnesha):

thanks.

OpenStudy (marziajackson$):

is it B?

OpenStudy (marziajackson$):

@Michele_Laino

OpenStudy (marziajackson$):

@Nnesha